# Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization

### v2

@@ 4.1.1 Protocol for Quantitative Measurement: a Hypothesis Testing Method
To distinguish the past and the future, we take the border as the release date of the model. For example, if a LLM is released at Jan. 2024 (also denoted as ‘cutoff time’ in this work), the time after Jan. 2024 would be the future. We could collect question prediction tasks in different periods. Based on different periods, We compute the prediction accuracy in these different periods, denoted as $acc_{\text{past}}$ and $acc_{\text{future}}$, where accuracy of the present (somewhere between the past and future) the is denoted as $acc_{\text{present}}$.
@@
$\small T_{1}:\ H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}],\$ (4)
@@
$\small T_{2}:\ H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}],$ (5)
@@
$\small\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
@@
Experimental setting We have sourced all questions from GoodJudgmentOpen [25] through web scraping and API downloads. The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing. Throughout this experiment, approximately 50 GPU hours on NVIDIA A800 were consumed.
We collected 2532 questions from GoodJudgmentOpen  [25], earliest one was proposed on 2015-09-01 and the most recent one was closed on 2024-05-25. It is "Will the adopted EU Artificial Intelligence (AI) Act be published in the Official Journal of the European Union before 25 May 2024?" Most common tags are in Appendix F.
@@ 4.1.2 Observation
Neophilia Bias is prevalent closer to release, whereas Nostalgia Bias occurs earlier. As shown in Tab. 2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent.
Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release. Pythia-12B and Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend. OPT-2.7B is Neophilia overall, but this observation is likely due to its consistently near-random performance across various periods, rather than a true preference for novelty.
@@ 4.2.1 Protocol
In the prospective view when especially using LLMs for prognostication prediction, there might be a temporal degeneration issue defined as below:
@@
Temporal Degeneration The capability of prognostication prediction might be degraded over time due to the larger generalization gap for longer time.
To qualitatively measure Temporal Degeneration, we propose a new hypothesis testing.
@@
$\small T_{3}:$ $\ H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}],$ (7)
@@
Table 3: Accuracy performance of various models across different periods. The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for ††, and p < 0.001 for ††*), benchmarked against their accuracy before release. Grey cells indicate the models have not been released in this period and represent the average scores of the models prior to their release.
Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4
OPT-13B 0.55 0.48 0.31** 0.39 0.41 0.52 0.37* 0.27**
OPT-2.7B 0.17 0.20 0.17* 0.22 0.25 0.22 0.26 0.26
LLaMA-7B 0.21* 0.17** 0.16** 0.22* 0.25 0.22** 0.22* 0.24*
Pythia-12B 0.48 0.47 0.32** 0.43 0.48 0.39* 0.43 0.31**
Falcon-rw-1B 0.22 0.16 0.14* 0.28 0.27 0.21 0.20 0.24
Baichuan-7B-Chat 0.32 0.24* 0.30 0.23 0.19** 0.20* 0.34
InternLM-Chat-7B 0.35 0.23** 0.33 0.27 0.29 0.30 0.31
GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.30**
GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.37** 0.48 0.40**
LLaMA-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40*
LLaMA-2-7B 0.27 0.26 0.27 0.22 0.26 0.34
LLaMA-2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38
Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.41
Baichuan2-13B-Base 0.49 0.30** 0.46 0.42 0.28** 0.40*
Baichuan2-13B-Chat 0.6 0.33** 0.43 0.41* 0.26** 0.27**
Zhongjing-Base 0.38 0.30 0.33 0.26* 0.39
Mistral-7B-v0.1 0.44 0.41 0.34* 0.28* 0.34*
Phi-1.5 0.41 0.38 0.39 0.30 0.25**
Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33*
Qwen-14B-Chat 0.39 0.25** 0.23** 0.33 0.27**
Qwen-7B 0.36 0.36 0.30 0.33 0.25**
Skywork-13B-Base 0.37 0.39 0.29* 0.30 0.27*
ChatGLM3-6B 0.42 0.38 0.33* 0.26** 0.34
Zephyr-7B-beta 0.4 0.32 0.31* 0.28* 0.28*
Yi-6B 0.29 0.27 0.16** 0.22 0.34
Yi-6B-Chat 0.38 0.25** 0.39 0.35
Qwen-1.8B 0.35 0.22** 0.37 0.30
RWKV-v5-Eagle-7B 0.29 0.29 0.33 0.28
GPT-4-231106 0.66 0.44** 0.39** 0.36**
Phi-2 0.31 0.23* 0.22 0.30
Command R+ 0.54 0.38**
Mixtral-8x22B-Instruct-v0.1 0.56 0.40**
Phi-3-mini-4k-instruct 0.38 0.26**
Qwen1.5-110B-Chat 0.56 0.40**
@@ 4.2.2 Observation
Earlier models perform worse. As the latest scores in Table 3 shows, models released earlier perform worse. The poor performance may be due to two factors. One is the natural decay over time, as evidenced by the clear downturn in LLaMA-7B’s performance. Over time, earlier models fail to adapt to new data. The other is the initial lack of advanced training techniques, which may have resulted in models like OPT showing near-random performance from the outset, without significant changes. This is mainly because earlier models were trained with less advanced techniques, leading to poorer performance.
Most models exhibit weak temporal generalization ability. Experimental results show that many models have significant fluctuations in accuracy percentages after release, indicating weak temporal generalization ability. Models with strong predictive capabilities fail to maintain their scores, even if only for a short period. For instance, GPT-4-230613 has demonstrated a significant decline in four out of six periods after release, while GPT-4-231106 has seen a significant decrease across all three observed periods. This variability and instability suggest that models have insufficient adaptive capabilities when dealing with new, unseen data. Some models exhibit near-random performance after the cutoff event, highlighting their limited understanding of the world and inability to generalize effectively from past information to future contexts. This shows that current models’ temporal generalization ability is significantly lacking and needs further improvement.
More advanced models perform better in future predictions. . For instance, Qwen1.5-110B-Chat, Command R+(104B), Mixtral-8x22B-Instruct-v0.1, along with GPT-4, have achieved the highest scores during the period from March to April 2024. It is worth noting that certain versions of LLaMA-2 and Baichuan2 have also performed commendably.
This observation may stem from a bias, as models like Qwen1.5-110B-Chat , Command R+(104B), Mixtral-8x22B-Instruct-v0.1 are the most recently released model, and a more robust evaluation awaits further data updates. The advanced training techniques employed by these models likely contribute to their enhanced accuracy and stability. As time advances and an increasing amount of data is accumulated, the models’ assessments are expected to become more precise and valuable.
5 Conclusion
@@
We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts. Our experiments reveal that LLMs often struggle with temporal biases, affecting reliability in dynamic scenarios requiring accurate forecasting. The evaluation framework and datasets we introduce allow for an in-depth assessment of model performance over time. The findings highlight shortcomings in existing benchmarks related to temporal shifts and underscore the urgent need for improvements in LLM training and updating processes to enhance adaptability and mitigate biases.
Limitations
While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language. Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
@@
Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
@@
Community Call to Action
We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events. This observation supports the findings in  [26], which suggest that models are most familiar with knowledge up to 2020. This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning. It suggests that caution is needed, especially for those hoping to use model judgments directly.
@@ 5.3 Wikipedia
Acknowledgement
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
Appendix A Testing Statistical Hypothesis

### v3

@@ 4.1.1 Protocol for Quantitative Measurement: a Hypothesis Testing Method
We discussed generally what is past, present, and future for a model in Sec 2.1.1. Now, let’s clarify the context of the prediction problem. For each prediction task, there is an initiation time (when the question is posed) and a close time (when the answer is revealed). If a question’s closure time precedes the model’s release date (for instance, by more than three months), then that question is considered part of the past for that model. The present refers to questions whose close time is near the release time of the model, capturing the period around when the model was launched.
@@
$T_{1}:$ $H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}],\quad$ (4)
$H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}]$
@@
$T_{2}:$ $H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}],\quad$ (5)
$H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}]$
@@
$\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
@@
Detailed experiment settings are listed in Appendix D.
@@ 4.1.2 Observation
Nostalgia Bias is prevalent.As shown in Tab. 2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent. Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release. Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend.
@@ 4.2.1 Protocol
We discussed generally Temporal Degeneration in 2.3, here we need to define specifically.
@@
Temporal Degeneration The capability of future event prediction might be degraded over time due to the larger generalization gap brought by time.
To qualitatively measure Temporal Degeneration, we propose a hypothesis testing like 4.1.1.
@@
$T_{3}:$ $H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}],$ (7)
$H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}]$
@@
Table 3: Accuracy performance of various models across different periods. The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for **, and p < 0.001 for ***), benchmarked against their accuracy before release. Grey cells indicate the models have not been released in this period and represent the average scores of the models prior to their release.
Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4 24/5 -24/6 24/7 -24/8
Baichuan-13B-Chat 0.52 0.26** 0.20** 0.36** 0.32** 0.33** 0.20** 0.23** 0.17**
GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.29** 0.28** 0.26**
GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.38** 0.48 0.40** 0.35** 0.36**
Llama-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40* 0.28** 0.34**
Llama-2-7B 0.27 0.26 0.27 0.22 0.26 0.33 0.18** 0.24
LLaMA2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38 0.27** 0.28**
Baichuan2-13B-Base 0.49 0.30** 0.46 0.42* 0.28** 0.39* 0.27** 0.27**
Baichuan2-13B-Chat 0.5 0.33** 0.43 0.41* 0.26** 0.29** 0.29** 0.33**
Baichuan2-7B-Base 0.42 0.26** 0.30* 0.24** 0.33 0.33* 0.25** 0.22**
Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.40 0.23** 0.27*
Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33* 0.25** 0.26**
Mistral-7B-v0.1 0.44 0.41 0.35* 0.28* 0.34* 0.25** 0.24**
Phi-1.5 0.41 0.38 0.38 0.30 0.26** 0.32* 0.27**
Qwen-14B-Chat 0.39 0.25** 0.24** 0.33 0.27** 0.28** 0.19**
Zephyr-7B-beta 0.4 0.32 0.32* 0.28* 0.28** 0.31* 0.19**
Yi-6B 0.29 0.27 0.16** 0.22 0.33 0.21* 0.21*
Gemini 0.43 0.20** 0.37 0.24** 0.21** 0.17**
Qwen-1.8B 0.35 0.22** 0.37 0.29 0.20** 0.20**
GPT-4-231106 0.66 0.45** 0.39** 0.36** 0.47** 0.43**
Phi-2 0.31 0.23* 0.22 0.29 0.18** 0.15**
Claude-3-opus-20240229 0.53 0.52 0.37** 0.47 0.36**
Claude-3-sonnet-20240229 0.23 0.24 0.13** 0.10** 0.07**
Command R+ 0.54 0.37** 0.43* 0.33**
DeepSeek-V2-Chat 0.58 0.49** 0.46** 0.35**
Mixtral-8x22B-Instruct-v0.1 0.56 0.39** 0.43** 0.35**
Phi-3-mini-4k-Instruct 0.38 0.26** 0.24** 0.26**
Qwen1.5-110B-Chat 0.56 0.40** 0.45* 0.36**
Claude-3.5-Sonnet-20240620 0.63 0.48** 0.23**
Gemini-1.5-Pro 0.5 0.47 0.35**
Gemini-1.5-flash 0.39 0.23** 0.22*
Qwen2-72B-Instruct 0.6 0.45** 0.40**
LLaMA-3.1-405B-Instruct 0.5 0.41
GPT-4o 0.62 0.39**
GPT-4o-mini-2024-07-18 0.48 0.28**
@@ 4.2.2 Observation
Earlier models perform worse, while newer advanced models perform better in future predictions. As the latest scores in Table LABEL:grey_table show, earlier models, show a significant performance decline across multiple time periods. For example, LLaMA-7B’s accuracy dropped from 0.11 in March 2023 to 0.08 by March 2024, with no improvement in later evaluation periods. In contrast, newer models exhibit much stronger future prediction capabilities.
Close-source models perform well initially but struggle with temporal generalization, while open-source models show better stability. As is shown in Appendix C, in both pre- and post-release scenarios, proprietary models such as GPT-4 and Claude-3.5 slightly outperform their open-source counterparts. However, when predicting outcomes for events known only post-release, all models exhibited substantial declines in performance.
Our analysis highlights models with less than 31% decline in post-release accuracy in bold, illustrating relatively stable performance. Conversely, models that experienced declines greater than 39% are marked by underline, indicating substantial performance degeneration. Notably, models such as Command R+, Mixtral-8x22B-Instruct-v0.1, DeepSeek-V2-Chat, and LLaMA-3.1-405B-Instruct demonstrate slower declines in accuracy post-release. Conversely, models like Claude-3.5-Sonnet-20240620, GPT-4-231106, and GPT-4o show rapid declines in performance,
The case studies of model iterations within the Claude series and the GPT models illustrate the dynamics of model evolution and generalization capabilities over time. Notably, the Claude-3-opus-20240229 model shows the least decline, indicating a potentially better tuning for broader applications or general datasets. This contrasts with the more significant declines seen in Claude-3-sonnet-20240229 and Claude-3.5-Sonnet-20240620, suggesting that these iterations may have undergone specific architectural tweaks, utilized different training data scopes, or implemented optimization strategies that did not generalize as effectively.
Similarly, the progression from GPT-3.5-turbo-230613 to GPT-4-231106 shows an increasing trend in both pre-release accuracy and the magnitude of decline post-release, with the GPT-4-231106 model experiencing the largest drop (39.39%). This trend could indicate a push towards optimizing these models for higher initial accuracy, possibly at the expense of their ability to generalize effectively over time.
Miniature versions of models show significant declines in temporal performance. The data reveals that miniature versions of models underperform compared to their full-sized counterparts in temporal tasks. For instance, Gemini.pro had an accuracy of 0.37 in March 2024, but this dropped sharply to 0.17 by July 2024, indicating a significant loss in precision for temporal tasks. Similarly, the mini version of GPT-4 showed considerably lower performance across multiple time periods, particularly in April 2024, where its accuracy was just 0.21, around 0.1 lower than the full-sized GPT-4 in the same period. This illustrates that while mini models improve inference efficiency, they exhibit poor generalization over time, with performance rapidly declining in long-term prediction tasks.
5 Conclusion and Community Call to Action
@@
We rigorously define and quantify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts. Our experiments reveal that LLMs often struggle with temporal biases, affecting reliability in dynamic scenarios requiring accurate forecasting. The evaluation framework and datasets we introduce allow for an in-depth assessment of model performance over time. The findings highlight shortcomings in existing benchmarks related to temporal shifts and underscore the urgent need for improvements in LLM training and updating processes to enhance adaptability and mitigate biases.
Specifically, we urge the community to take note of this phenomenon: models tend to better comprehend earlier events.  Cheng et al. (2024) suggests that models are most familiar with knowledge up to the year 2020. However, our findings indicate that the understanding of world events (even in cases where models can remember) lingers around 2015, with a trend towards even earlier years (beyond the time range of our test data). This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning. It suggests that caution is needed, especially for those hoping to use model judgments directly.
Acknowledgements
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
@@ Acknowledgements
Looking ahead, we suggest future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
  • Exploring mitigation strategies, through we conducted preliminary investigations into prompt engineering strategies that could potentially mitigate temporal performance degradation, see Abstract B.
@@
Limitations
While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language. Also, the English-centric nature of the questions and answers limits the benchmark’s applicability to non-English speaking regions and global events. Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
Potential Risks
Bias and Fairness Risks: When used to make predictions about future events, language models can inadvertently perpetuate or amplify biases present in their training data. This could lead to biased forecasts that reflect and potentially exacerbate existing societal prejudices. Such biased outputs may result in models that are unfair or discriminatory, which could harm individuals or groups and damage the reputation of the organizations involved.
@@
Appendix A Related Work
Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively. Recently, several methods have been proposed to evaluate models at different times and in various contexts Fatemi et al. (2024); Kasai et al. (2024); Vu et al. (2023). TimeQA Chen et al. (2021) evaluates a model’s ability to handle temporal evolution, while Situated QA Zhang and Choi (2021) focuses on a model’s performance in specific contexts. StreamingQA Liška et al. (2022) introduces the concept of continuous information flow, while RealTimeQA Kasai et al. (2024) and FreshQA Vu et al. (2023) emphasize real-time and up-to-date information processing capabilities. LatestEval  Li et al. (2023b) utilizes the most recent text to construct questions and answers, covering six most frequently queried categories including explanation, summary, reason, demonstrations, existence, and possible usage.
However, it’s essential to clarify that RealTimeQA Kasai et al. (2024) and FreshQA Vu et al. (2023), which continuously update topics, are inclined towards knowledge. For example, RealTimeQA Kasai et al. (2024) has such a sample in their paper, "Which wildly popular show was recently greenlit for a new season?" (Answer: Squid Game). These methods are typically used for evaluating retrieval-augmented generation (rag) series models. On the other hand, methods like LatestEval  Li et al. (2023b), which construct questions using the most recent information, lean towards enhancing reading comprehension.
In contrast to these approaches, our focus is on testing the model’s understanding of the world, as demonstrated through prediction. Our method of evaluation is based on facts and computation loss. This ensures that our evaluation is resistant to data contamination, objective, and free from biases in question construction and judgment. This approach allows us to examine the model’s ability to process and predict real-world events accurately and efficiently.
Future-knowledge Prediction
Previous research has focused on how to automatically predict real-world events. For instance, Zou et al. (2022) constructed a dataset of questions obtained from tournament predictions as an evaluation benchmark for automatic forecasting. However, this dataset may already be outdated because its knowledge is likely already embedded within current LLMs. Recently, studies have shown that LLMs can match human performance in the task of forecasting the future. For example, Schoenegger et al. (2024) demonstrates that the predictive performance of an ensemble of multiple LLMs is superior to that of a single LLM, and it even surpasses the performance of a group of human forecasters.
Model selection and Experimental Setting We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems. They are not only benchmarks of their respective sizes, ranging from 1 billion to 40 billion parameters, but also exemplify the rapid progression in NLP technology. We set the context size to 2048, then compute the sum of the negative logarithms of each segment. We then divide the computed sum by the length of the text under UTF-8 encoding. Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
Appendix B Exploring Mitigation Strategies
While our primary focus is temporal evaluation rather than solution development, we conducted preliminary investigations into prompt engineering strategies that could potentially mitigate temporal performance degradation.
B.1 Try Different Prompt to Mitigate
Motivated by observed safety constraints and refusal patterns in base models and the result of Pham and Cunningham (2024), we designed five specialized prompts to investigate how different framing strategies affect temporal prediction:
Figure 6: Accuracy of LlaMA3.1-8B-Instruct with different prompts
Figure 7: Accuracy of Qwen2.5-7B with different prompts
Figure 8: Accuracy of Baichuan2-7B-Chat with different prompts
Figure 9: Accuracy of Yi-6B with different prompts
  • Base: Used direct incentivization ("tip of \$200") and explicit guessing encouragement to counter safety-driven refusals.
  • Dream Framework (v1): Encourages speculative thinking by positioning the model as a visionary oracle.
  • Time Travel (v2): Frames predictions as historical memories from a future perspective.
  • Multiverse (v3): Leverages parallel universe observations to bypass temporal constraints.
  • Precise Thinking (v4): Emphasizes structured temporal reasoning chains.
  • Future Recall (v5): Simulates retrospective analysis from a future vantage point.
See Figure 9, 9, 9 and 9. Specifically, different models respond variably to the designed prompts. For Yi-6B, there were scarcely any discernible differences in performance across the several prompts we employed. In contrast, for the Qwen2.5-7B model, all modified prompts generally outperformed the base prompt in most scenarios. Analyzing the results from Llama3.1-8B-Instruct, Qwen2.5-7B and Baichuan2-7B-Chat, we found that prompts v5(Future Recall) and v4(Precise Thinking) demonstrated relatively better effectiveness. The specific impacts vary from model to model, which is closely related to the model’s instruction-following capabilities.
In addition, our exploratory analysis revealed significant performance variations across question categories, with the model achieving lowest accuracy on "Finance" (58.1%), "Business" (55.4%). and "Economic Indicators" (62.3%) categories, compared to stronger performance on "Foreign Policy", (81.2%), "Security and Conflict" (79.6%) and "Leader Entry/Exit" (83.7%) categories.
It is plausible that fine-tuning models with analysis and prediction data (especially in these weak areas) could fully unlock their inner predictive potential.
Appendix C Full Table of Decline in Prediction Accuracy
See Table 4.
Table 4: Comparison of Powerful Open-source and Closed-source LLMs
Model Pre-release Accuracy First Post-release Accuracy Decline Percentage Decline
Baichuan-13B-Chat 0.52 0.23 0.29 55.77%
GPT-3.5-turbo-230613 0.49 0.33 0.16 32.65%
GPT-4-230613 0.60 0.41 0.19 31.67%
Llama-2-13B 0.52 0.45 0.07 12.50%
Llama-2-7B 0.27 0.27 0.01 1.85%
LLaMA2-7B-Chat 0.45 0.38 0.07 15.56%
Baichuan2-13B-Base 0.49 0.38 0.11 22.45%
Baichuan2-13B-Chat 0.50 0.38 0.12 24.00%
Baichuan2-7B-Base 0.42 0.28 0.14 33.33%
Baichuan2-7B-Chat 0.37 0.39 -0.02 -5.41%
Colossal-LLaMA-2-7B-Base 0.42 0.40 0.02 4.76%
Mistral-7B-v0.1 0.44 0.38 0.06 13.64%
Phi-1.5 0.41 0.38 0.03 7.32%
Qwen-14B-Chat 0.39 0.24 0.15 37.18%
Zephyr-7B-beta 0.40 0.32 0.08 20.00%
Gemini 0.43 0.29 0.14 33.72%
Yi-6B 0.29 0.22 0.07 25.86%
Qwen-1.8B 0.35 0.29 0.05 15.71%
GPT-4-231106 0.66 0.42 0.24 36.36%
Phi-2 0.31 0.23 0.08 27.42%
Command R+ 0.54 0.40 0.14 25.93%
Claude-3-opus-20240229 0.53 0.45 0.09 16.04%
Claude-3-sonnet-20240229 0.23 0.18 0.05 19.57%
DeepSeek-V2-Chat 0.58 0.47 0.10 18.10%
Mixtral-8x22B-Instruct-v0.1 0.56 0.41 0.15 26.79%
Phi-3-mini-4k-Instruct 0.38 0.25 0.13 34.21%
Qwen1.5-110B-Chat 0.56 0.43 0.14 24.11%
Claude-3.5-Sonnet-20240620 0.63 0.35 0.28 43.65%
Gemini-1.5-Pro 0.50 0.41 0.09 18.00%
Gemini-1.5-flash 0.39 0.23 0.17 42.31%
Qwen2-72B-Instruct 0.60 0.43 0.17 29.17%
LLaMA-3.1-405B-Instruct 0.50 0.41 0.09 18.00%
GPT-4o 0.62 0.39 0.23 37.10%
GPT-4o-mini-2024-07-18 0.48 0.28 0.20 41.67%
Appendix D Experimental setting
For this experiment, we sourced all questions from Good Judgment Open using web scraping. The data collected includes initiation and conclusion dates, domains, questions, descriptions, and answers, which were appropriately ordered for further processing. Approximately 50 GPU hours on NVIDIA A800 were consumed during this experiment.
Up to 2024.10.1, we collected a total of 2769 questions from Good Judgment Open . The earliest question was proposed on 2015-09-01: "Will Congress pass a resolution disapproving the Joint Comprehensive Plan of Action?" The most recent question was closed on 2024-10-01: "What will be the 12-month percentage change in the US Consumer Price Index (CPI) for September 2024?" Question samples and most common tags are in Appendix J.
Appendix E Testing Statistical Hypothesis